The battery powers 4 motors in the drone. 

Each motor has a resistance of 1.60 Ω when the power input to each motor is 19.6 W 

The 4 motors are connected in parallel with the battery. 

Calculate the current through the battery. 


1
Expert's answer
2021-02-03T16:17:20-0500

The current through each motor:

"I_1=\sqrt{\frac{P}{R}}=\sqrt{19.6/1.6}=3.5A"

The current through the battery:

"I=4I_1=4\cdot3.5=14A"